# Preprint Manuscript

### v3

@@
Preprint Manuscript
Third Version
CUBIC INCOMPLETENESS: HILBERT’S TENTH PROBLEM OVER $\mathbb{N}$ STARTS AT $\delta=3$
Milan Rosko
(Oktober 2025)
We prove that Hilbert’s Tenth Problem over $\mathbb{N}$ remains undecidable when restricted to cubic equations (degree $\leq 3$), resolving the open case $\delta=3$ identified by Jones (1982) and establishing sharpness against the decidability barrier at $\delta=2$ (Lagrange’s four-square theorem). For any consistent, recursively axiomatizable theory $T$ with Gödel sentence $G_{T}$, we effectively construct a single polynomial $P(x_{1},\ldots,x_{m})\in\mathbb{Z}[\mathbf{x}]$ of degree $\leq 3$ such that $T\vdash G_{T}$ if and only if $\exists\mathbf{x}\in\mathbb{N}^{m}:P(\mathbf{x})=0$. Our reduction proceeds through four stages with explicit degree and variable accounting. Proof-sequence encoding via Diophantine $\beta$-function and Zeckendorf representation yields $O(KN)$ quadratic constraints ($K=O(\log(\max_{i}f_{i}))$, NN = proof length). Guard-gadgets wrap each constraint $E(\mathbf{x})=0$ as $u\cdot E(\mathbf{x})=0$, $u=1+v^{2}$, preserving degree $\leq 3$ with $O(KN^{3})$ variables. Sum-of-squares aggregation $P_{\text{merged}}=\sum_{i}P_{i}^{2}$ produces degree $\leq 6$ with $O(KN^{3})$ monomials. Recursive monomial shielding reduces degree to $\leq 3$ in $O(\log d)$ rounds, adding $O(K^{3}N^{3})$ variables total.
Keywords: Proof Theory, Logic, Hilbert’s Tenth Problem, Zeckendorf’s Theorem
Open Problem, Constructive Mathematics, Fibonacci, Cubic equations
Primary MSC: 11U05, 03F40, 03D35
Secondary MSC: 11D72, 03B25, 11Y16
Milan Rosko is from University of Hagen, Germany
Email: Q1012878 @@ studium.fernuni-hagen.de – ORCID: 0009-0003-1363-7158
Licensed under “Deed” ΓΔ  – creativecommons.org/licenses/by/4.0
1 Corrigendum
1.1 First Corrigendum
The initial version encoded a Gödel sentence as a Diophantine system rather than a single equation. The ambition to present a concrete standalone equation introduced an unwarranted bias into the construction. Subsequent algorithmic approaches yielded a massive closed form, but the reduction diverted attention from the central epistemic questions.
1.2 Second Corrigendum
In the second revision, we pivoted to explicit witness construction, ensuring all constraints remain genuinely cubic, and grounded the argument in logical foundations rather than forcing a premature encoding.
1.3 Third Corrigendum
In this third—still preliminary—version, we resolve ambiguities between logical and mathematical layers ($\mathbb{Z}$ versus $\mathbb{N}$), and improve on bookkeeping. The first-order argument is developed independently of its Diophantine realization. We provide a high-level overview at the outset, then verify each layer on its own terms.
@@ 1.3 Third Corrigendum
We now adopt a direct approach: all variables, constraints, and solutions are understood to inhabit $\mathbb{N}$ from the outset. The $\beta$-function (Theorem 3.1), four-squares encoding (Lemma 4.2), and Zeckendorf representation (Definition 3.3) operate naturally over $\mathbb{N}$. The guard-gadget framework (Definition 3.6) enforces positivity via $u=1+v^{2}\geq 1$ without recourse to negative integers. This interpretation aligns with the historical formulation of Hilbert’s Tenth Problem over $\mathbb{N}$ and avoids unnecessary detours through $\mathbb{Z}$. The approach is straightforward: Turing machines operate with discrete symbols.
1.4 Context
This paper shares ideas with different work [^1] The Fractal Logic of $\Phi$-adic Recursion, arXiv:2510.08934, [2025a], in that a compression-undecidability relationship is visible in the context of extensions from $\mathbb{N}$ to $\mathbb{Z}$. The additional compression offered by negative numbers permits lower polynomial degrees in Diophantine equations, but at the cost of increased uncertainty between powers. The equation $x^{2}=4$ is uniquely solved in $\mathbb{N}$ by $x=2$, while in $\mathbb{Z}$ the solution set comprises $\{-2,2\}$. More dramatically, $x^{4}=16$ admits four solutions $\{\pm 2,\pm 2i\}$ in $\mathbb{C}$, with the real solutions $\pm 2$ already exhibiting the doubling. Future work will focus on this: The ambiguity is not a deficiency but a necessary consequence of Kolmogorov [1968] complexity: syntactic reduction of degrees is purchased through semantic branching. Results possessing nontrivial structure in $\mathbb{N}$ collapse in $\mathbb{Z}$ to trivial symmetries, whereby the displaced complexity migrates into sign ambiguity.
2 Overview
2.1 The Problem
Previously [^2] A Fibonacci-Based Gödel Numbering: $\Delta_{0}$ Semantics Without Exponentiation, arXiv:2509.10382, [2025b], we proposed that a Zeckendorf-based arithmetization of Gödel numbering—eschewing exponentiation in favor of additive structure—might lower the formal overhead in encoding undecidability. We now apply this method to produce undecidable Diophantine equations of degree three, thereby advancing “Hilbert’s Tenth Problem” [^3] As posed by Hilbert [1900]. into the cubic regime ($\delta=3$) and resolving a previously open case: We cite from the article “Universal Diophantine Equation”, by Jones [1982]:
Hence, we define the open question:
Problem 2.1 (Cubic Remainder of Hilbert’s 10th Problem).
Let $\delta=3$ denote the set of all polynomial Diophantine equations of the form
$P(x_{1},\ldots,x_{n})=0,\quad P\in\mathbb{N}[x_{1},\ldots,x_{n}],\quad\delta(P)\leq 3.$ (2.1)
Question: Is there an algorithm that, given $P$ and $\delta$ = 3, decides whether
2.2 Roadmap

### v4

# CUBIC INCOMPLETENESS: HILBERT’S TENTH PROBLEM AT DEGREE THREE

@@
Abstract
We analyze the cubic fragment $\mathcal{D}_{3}$ over $\mathbb{N}$ by isolating the uniform closure principle any total correct cubic solver would have to realize. In $\mathsf{HA}$ we give a fully constructive, additive and degree-controlled encoding of bounded verification: for each externally fixed bound, we effectively produce a finite system of degree-3 Diophantine equations whose solutions represent the existence of the corresponding finite proof or computation trace. The encoding is purely syntactic, using “gadgets” and Carryless Pairing. In a classical metatheory (e.g. $\mathsf{PA}$) we show that the global solver hypothesis implies a uniform operator eliminating the bound inside $\mathcal{D}_{3}$, which is incompatible with standard non-uniformity/realizability constraints. Hence no uniform cubic can exist clasically.
Subject: Proof Theory, Diophantine Geometry, Computability Theory
1 Introduction
Exposition.
Figure 1: The construction proceeds through strictly separated layers. Undecidability is established at the logical level; a classical reflection step then shows that no algorithm can solve the problem uniformly, due to an intrinsic global barrier. The proof is stratified across constructive, equational, and reflective layers.
Previously [2025], we proposed that a Zeckendorf-based arithmetization of Gödel numbering—eschewing exponentiation in favor of additive structure—can lower the formal overhead of undecidability encodings. We now apply this method to the cubic regime, addressing the degree-$3$ remainder singled out by Jones [1982] in Universal Diophantine Equation:
The guiding analogy is the classical impossibility of the Delian problem, best read as a closure failure. Straightedge–compass constructions generate exactly the constructible numbers: the smallest class containing $\mathbb{Q}$ and closed under the permitted operations (equivalently, under towers of quadratic extensions). Doubling the cube would require closing these rules under a genuinely new operation (taking cube roots). The obstruction is therefore global: no amount of local ingenuity can replace a missing closure principle.
Our degree-$3$ analysis has the same logical shape (though the closure is logical rather than algebraic). A uniform solver hypothesis for $\mathcal{D}_{3}$ would have to supply an additional closure step: a uniform elimination of the external bound inside the cubic fragment (a cubic closure operator). The decisive barrier is thus not an algebraic accident (a particular gadget failing), but the absence of a global closure principle: bounded representability holds, while uniform existential closure is the forbidden move.
2 Ansatz
2.1 Core Phenomena
Problem 2.1 (Cubic remainder of Hilbert’s Tenth Problem over $\mathbb{N}$).
Let
$\mathcal{D}_{3}\;\equiv\;\{\,P\in\mathbb{Z}[x_{1},\ldots,x_{n}]:\delta(P)\leq 3\text{ for some }n\,\}.$ (2.1)
Question: Is there an algorithm which, given a code for $P\in\mathcal{D}_{3}$, decides whether
Observation.
We read the degree-$3$ remainder of Hilbert’s Tenth Problem as a closure problem rather than as a missing gadget or an incremental technical gap between the known degree-$4$ universal constructions and the degree-$2$ decidable regime. The decisive distinction is between:
  1. (i)
    bounded representability (local, finitary verification), and
  2. (ii)
    uniform existential closure.
At degree $3$ the cubic fragment is already expressive enough to represent, with externally fixed bounds, the finite witnesses that arise in verification phenomena: bounded computation traces, bounded proof traces, and other finitary tableaux. Formally, for broad families of bounded predicates $R_{B}(x)$ (with $B$ an external numerical bound), one can effectively associate cubic Diophantine instances $P_{x,B}\in\mathcal{D}_{3}$ such that
$R_{B}(x)\quad\Longleftrightarrow\quad\mathsf{Sol}_{3}(\ulcorner P_{x,B}\urcorner),$ (2.3)
uniformly in $(x,B)$. The uniform decision problem, however, is inherently unbounded: it asks for a single algorithm deciding the existential closure
$\exists B\;R_{B}(x).$ (2.4)
Any putative uniform solver for $\mathcal{D}_{3}$ would therefore have to implement, in one form or another, a uniform closure principle that internalizes this unbounding step within the cubic fragment: a uniform procedure that, given $x$, compiles the family $(P_{x,B})_{B\in\mathbb{N}}$ into a single cubic instance $Q_{x}\in\mathcal{D}_{3}$ satisfying
$\exists B\;R_{B}(x)\quad\Longleftrightarrow\quad\mathsf{Sol}_{3}(\ulcorner Q_{x}\urcorner).$ (2.5)
This is not an auxiliary convenience but the logical core of uniformity: a uniform solver is, in effect, precisely a uniform method for closing the external existential bound. In this sense the cubic remainder is self-similar. Degree $3$ is the first regime in which bounded verification can be arithmetized cleanly while the fragment still fails to be closed under the one semantic operation that would turn bounded representability into a global decision procedure. Degree $4$ has enough algebraic slack to absorb the closure move (hence universality), degree $2$ lacks the expressive room to host the bounded encodings at all, and degree $3$ is the critical boundary.
2.2 Logical Architecture and Scope
Exposition.
We analyze the proof-theoretic content of a global solver hypothesis for the cubic fragment and isolate the uniform closure principle that such a hypothesis would have to validate under a constructive reading. The contradiction is obtained by a stratified Modus Tollens argument: we (i) formulate the solver hypothesis as a single arithmetical commitment, (ii) identify the uniform witness-producing content this commitment would carry under the intended constructive semantics, and (iii) refute precisely that witness-producing content by a realizability/non-uniformity barrier. A final reflection step is used only to pass from the constructive impossibility of realizers to the classical negation of the corresponding existence claim.
Ansatz 2.2.
Hilbert’s Tenth Problem, cf. Hilbert [1900], is a uniformity question: it asks whether there exists a single effective procedure which, given an arbitrary Diophantine equation from a prescribed class, decides solvability over $\mathbb{N}$. In this paper the class of interest is the cubic fragment
$\mathcal{D}_{3}\;\equiv\;\{\,P\in\mathbb{Z}[x_{1},\ldots,x_{n}]:\delta(P)\leq 3\text{ for some }n\,\}.$
The corresponding global statement is therefore:
There is no uniform algorithm which, on input $P\in\mathcal{D}_{3}$, decides whether $\exists\vec{x}\in\mathbb{N}^{n}\;P(\vec{x})=0$.
Importantly, we do not pursue this by exhibiting a distinguished cubic polynomial whose solvability is independent of a background theory, nor by constructing a universal cubic polynomial in the classical MRDP sense, cf. Matiyasevich [1993]. The obstruction analyzed here is global and logical: it is the incompatibility of a uniform method with the constructive non-uniformity phenomena made explicit below.
Definition 2.3 (Cubic Solvability Predicate).
Fix a Gödel Numbering $p\mapsto P_{p}$ of integer-coefficient polynomials together with their arities. Write $p\in\mathrm{Code}(\mathcal{D}_{3})$ if $p$ codes a polynomial of total degree at most $3$. Define the arithmetical predicate
$\mathsf{Sol}_{3}(p)\;:\Longleftrightarrow\;\bigl(p\in\mathrm{Code}(\mathcal{D}_{3})\bigr)\ \wedge\ \exists\vec{x}\in\mathbb{N}^{n(p)}\;P_{p}(\vec{x})=0.$
Definition 2.4 (Total Correct Cubic Solver).
Let $\mathsf{Run}(e,p,t,b)$ be the standard primitive recursive predicate expressing:
“The machine with code $e$ halts on input $p$ within $t$ steps and outputs the bit $b\in\{0,1\}$.”
We formalize the claim
$e$ is a total correct solver for $\mathcal{D}_{3}$ over $\mathbb{N}$
as the closed arithmetical sentence:
$\mathsf{Solve}_{3}(e)\;\Longleftrightarrow\;\forall p\Bigl($ $\,p\in\mathrm{Code}(\mathcal{D}_{3})\ \to\ \exists t\,\exists b\in\{0,1\}\;$
$\bigl(\mathsf{Run}(e,p,t,b)\ \wedge\ (b=1\leftrightarrow\mathsf{Sol}_{3}(p))\bigr)\Bigr).$ (2.6)
Remark.
Equation (2.4) is the global reading of the solver hypothesis: it is a single $\forall\exists$-commitment asserting totality and correctness on all cubic inputs. It is not an instruction to perform an external search over $(t,b)$; rather, it asserts that for each input $p$ there exists a terminating computation trace witnessing the solver’s output and that the output is correct.
The analysis in this paper is intentionally proof-theoretic, cf. Boolos et al. [2007]. We do not assume that from the mere (classical) truth of $\mathsf{Solve}_{3}(e)$ one can extract witnesses. Instead, we study what follows under a constructive reading: in settings such as in Heyting Arithmetic ($\mathsf{HA}$) equipped with a sound realizability interpretation, cf. Troelstra and van Dalen [1988]; Heyting [1930] for the relevant $\forall\exists$-fragment, a derivation of (2.4) entails a uniform witness-producing method extracting $(t,b)$ (together with a correct trace) from $p$. This is the precise sense in which uniformity becomes logically explicit.
Ansatz 2.5 (Stratified Method).
The argument is deliberately stratified into three layers, each with its own role, see Figure 1. We keep the explicit encoding separate from its metatheoretic reading.
Layer I. Working in $\mathsf{HA}$, we build explicit cubic Diophantine gadgets for bounded verification phenomena: finite proof traces and finite computation traces. All constructions here are effective and degree-controlled, but the relevant bounds (proof length, runtime) are supplied externally.
Layer II. We apply equisatisfiable algebraic transformations that reorganize bounded systems without changing solvability over $\mathbb{N}$. The purpose of this layer is bookkeeping: it normalizes shape while preserving the degree bound and makes explicit where standard “packaging” moves attempt to leave the cubic regime.
Layer III. In a classical metatheory (one may take $\mathsf{PA}$ for definiteness), we treat the solver hypothesis as a definite arithmetical commitment. The only use of classical reasoning at this layer is a reflection step applied to program-existence claims singled out by the constructive analysis: once the relevant realizer is shown not to exist, Modus Tollens yields the classical negation of the corresponding solver hypothesis.
Remark.
The paper isolates two independent non-uniformities and keeps them disjoint.
  1. (i)
    Bound selection. There is no computable function that, from $(\mathrm{T},\varphi)$, produces a bound sufficient for proof search. This is the classical incompleteness/undecidability non-uniformity.
  2. (ii)
    Cubic closure. Even when bounded verification is cubic-representable for each fixed bound, eliminating an unbounded existential parameter inside the cubic fragment is an additional uniformity demand. This closure demand is the crux of the cubic regime.
Our guiding Ansatz is that the degree-$3$ boundary is governed by the second barrier: at degree $3$ one can represent bounded verification cleanly, but one cannot uniformly close the external bound in a manner compatible with the intended constructive semantics.
Exposition.
Sections 34 implement Layer I: we give explicit cubic representations of bounded traces (proof checking and machine runs). Section 5.1 discusses Layer II normalization and makes precise where standard packaging steps request a closure move. The concluding section formulates the closure principle extracted from the solver hypothesis (under the explicit seam-check of the realizability interpretation) and performs the Layer III reflection step refuting $\exists e\,\mathsf{Solve}_{3}(e)$.
2.3 Bounded verification is cubic-representable
Exposition.
The role of Layer I is sharply delimited: we only represent bounded verification predicates. Given an external bound (proof length, runtime), brute-force simulation is decidable and yields a finite trace. Our technical task is to express the existence of such a finite trace by a Diophantine condition while keeping total degree $\leq 3$.
The point of doing this is not to “solve” an unbounded problem in disguise, but to isolate where the unbounded step must occur. Once the bounded predicate is represented cubically, any global solver for $\mathcal{D}_{3}$ that purports to decide unbounded solvability will necessarily have to internalize (i.e. existentially close) the bound. That closure is the crux addressed in the next subsection.
Definition 2.6 (Bounded Run Predicate).
Let $\mathsf{Run}(e,p,t,b)$ be the primitive recursive predicate:
“The machine with code $e$ halts on input $p$ within $t$ steps and outputs $b\in\{0,1\}$.”
We treat $(e,p,t,b)$ as external numerical parameters in the bounded verification problem.
Theorem 2.7 (Cubic Representability of Bounded Machine Runs).
There is a computable procedure which, given $(e,p,t,b)$, outputs a finite Diophantine system
$\Phi_{e,p,t,b}(\vec{z})=0$ (2.7)
over $\mathbb{N}$ such that:
  1. (i)
    every equation in $\Phi_{e,p,t,b}$ has total degree $\leq 3$, and
  2. (ii)
    $\exists\vec{z}\in\mathbb{N}^{m(e,p,t,b)}\;\Phi_{e,p,t,b}(\vec{z})=0\quad\Longleftrightarrow\quad\mathsf{Run}(e,p,t,b).$ (2.8)
Ansatz 2.8.
Fix $(e,p,t,b)$. A bounded computation can be written as a finite tableau (configurations at times $0,1,\ldots,t$) together with local transition constraints. We encode the tableau using low-overhead storage (e.g. $\beta$-style remainder storage for finite sequences) and enforce the transition relation by polynomial equalities whose arities are fixed and local.
Inequalities and boundedness constraints are converted to equations by standard sum-of-squares mechanisms. Degree inflation is avoided by a guard economy: only those constraints that require positivity are guarded, while already-activated or intrinsically bounded equations are left unguarded.
Remark.
Boundedness is essential. The system $\Phi_{e,p,t,b}$ is constructed with $t$ fixed externally. Thus Theorem 2.7 yields a family of cubic instances indexed by the bound. This is not yet a representation of $\exists t\,\mathsf{Run}(e,p,t,b)$, and no such unbounded closure is attempted in Layer I.
Definition 2.9.
Fix a recursively axiomatized theory $\mathrm{T}$ with a fixed proof system and Gödel Numbering. Write $\mathsf{Prf}_{\mathrm{T}}(\pi,\varphi)$ for the primitive recursive relation expressing:
$\pi$ codes a valid $\mathrm{T}$-proof of $\varphi$.”
Define the bounded provability predicate
$\mathsf{Prov}_{\mathrm{T},\leq N}(\varphi)\;\Longleftrightarrow\;\exists\pi\;\bigl(|\pi|\leq N\ \wedge\ \mathsf{Prf}_{\mathrm{T}}(\pi,\varphi)\bigr),$ (2.9)
where $N$ is treated as an external numerical parameter.
Theorem 2.10 (Cubic Representability of Bounded Proof Verification).
For every fixed $(\mathrm{T},\varphi,N)$ there is an effectively constructible finite Diophantine system
$\Psi_{\mathrm{T},\varphi,N}(\vec{x})=0$ (2.10)
over $\mathbb{N}$ such that:
  1. (i)
    every equation in $\Psi_{\mathrm{T},\varphi,N}$ has total degree $\leq 3$,
  2. (ii)
    and,
    $\exists\vec{x}\in\mathbb{N}^{m(\mathrm{T},\varphi,N)}\;\Psi_{\mathrm{T},\varphi,N}(\vec{x})=0\quad\Longleftrightarrow\quad\mathsf{Prov}_{\mathrm{T},\leq N}(\varphi).$ (2.11)
Exposition.
The system $\Psi_{\mathrm{T},\varphi,N}$ expresses the existence of a length-$N$ proof trace by:
  1. (i)
    storing the finite sequence of line-codes using $\beta$-style sequence storage,
  2. (ii)
    enforcing local verification constraints (axiomhood / inference steps) by low-degree equations controlled by Boolean activations, and
  3. (iii)
    using a carryless normal-form discipline for additive bookkeeping so that value-equalities can be imposed without explicit carry propagation gadgets.
All constraints remain within degree $3$ by construction.
Remark.
Theorem 2.10 is a bounded statement. It yields a family $(\Psi_{\mathrm{T},\varphi,N})_{N\in\mathbb{N}}$ of cubic instances. Layer I does not attempt to eliminate $N$ and does not claim a single cubic equation equivalent to $\mathrm{T}\vdash\varphi$. The passage from bounded verification to unbounded existence is exactly the closure step isolated in the next subsection.
Remark.
Throughout Layer I we maintain the following working invariant: constraints are guarded only when a positivity or nontrivial implication direction must be enforced. Equations that are already:
  1. (i)
    Boolean-bounded (e.g. b(b1)=0b(b-1)=0), or
  2. (ii)
    activated (multiplied) by a Boolean selector, or
  3. (iii)
    subordinate to a globally guarded block, are left unguarded.
This reduces notational overhead without changing satisfiability over $\mathbb{N}$ and preserves the degree bound.
2.4 Declarative Arithmetic
Exposition.
A recurring source of degree growth in Diophantine encodings is the temptation to implement arithmetic procedurally at the digit level. In standard base-$p$ representations, verifying an equality such as $x+y=z$ by digit-by-digit computation introduces explicit carry variables and normalization constraints; these quickly force higher-degree polynomials once one insists on uniformity.
Our approach is instead declarative. We enforce arithmetic equalities at the value level (e.g. $z=x+y$), and separately enforce that each integer appearing in the encoding has a canonical digit normal form subject only to local constraints. Uniqueness of normal forms then pins down the intended digits without any explicit carry propagation gadget. Existential quantification performs the “guessing”; the equations perform the verification.
Proposition 2.11 (Declarative Addition Principle).
Let $\mathcal{R}$ be a numeral system assigning to each $n\in\mathbb{N}$ a digit vector $\mathrm{dig}(n)$, such that:
  1. (i)
@@ 2.4 Declarative Arithmetic
  2. (ii)
    Uniqueness, each $n$ has a unique valid digit vector.
Then, for arithmetic constraints used inside a Diophantine encoding, it is not necessary to implement digit-level addition with carry gadgets. One may impose the value equality (e.g. $z=x+y$) and require valid digit vectors for $x,y,z$; uniqueness then forces $\mathrm{dig}(z)$ to coincide with the canonical normal form of $x+y$.
Remark.
Proposition 2.11 is not a claim about computability of addition in $\mathcal{R}$ (which may be expensive). It is a claim about verification: the solver supplies candidate digits existentially, and the constraint system certifies that these digits form a valid normal form for the integer already fixed by value-level equations.
Exposition.
We instantiate the above using Zeckendorf Representation. Write $F_{1}=F_{2}=1$ and $F_{k+2}=F_{k+1}+F_{k}$. Every $n>0$ has a unique representation
$n=\sum_{\kappa\in S}F_{\kappa},\quad S\subseteq\{2,3,\ldots\},\quad\text{with no consecutive indices.}$ (2.12)
Fix a cutoff $K$ (chosen externally in bounded settings), introduce digits $d_{2},\ldots,d_{K}\in\mathbb{N}$, and enforce:
$d_{\kappa}-d_{\kappa}^{2}=0,\quad d_{\kappa}d_{\kappa+1}=0,\quad n=\sum_{\kappa=2}^{K}F_{\kappa}d_{\kappa}.$ (2.13)
These are quadratic constraints. In particular, Zeckendorf Representation is verified without carry propagation.
Remark.
In base-$p$ systems, normal form is not stable under addition without explicit carry variables. By contrast, Zeckendorf Representation is enforced by a local forbidden-pattern condition (no adjacent $1$’s) plus reconstruction. When we impose a value-level equality $z=x+y$, the digits of $z$ are not computed; they are selected existentially and then uniquely determined by the normality constraints.
Exposition.
In Theorems 2.7 and 2.10 we repeatedly face constraints of the form
$u=u_{1}+u_{2},\quad u=u_{1}+u_{2}+u_{3},$ (2.14)
arising from local transition rules, pairing functions, or proof-rule bookkeeping. The naive approach would encode these equalities at the digit level (introducing carries) to maintain boundedness. We instead enforce the value equalities directly and require each relevant integer to carry a form in Zeckendorf Representation. By uniqueness, this pins down the digits without introducing global carry-propagation constraints and keeps the degree within the cubic regime once guards and activations are accounted for.
Remark.
In Layer I all digit systems have externally fixed width (cutoff $K$), because we only represent bounded traces. The dependence of $K$ on the size of the bounded objects is harmless there: $K$ is part of the externally supplied bound data. The unbounded elimination of such parameters is not attempted in Layer I; it is precisely the closure issue isolated next.
2.5 The cubic closure problem
Exposition.
Layer I yields cubic representations of bounded verification predicates. The global solver hypothesis, however, speaks about unbounded existence: whether there exists some witness (some runtime, some proof length, some tableau size) satisfying the bounded predicate.
Thus the decisive step is the passage from a family of bounded instances
$(\Phi_{e,p,t,b})_{t\in\mathbb{N}}\quad\text{representing}\quad\mathsf{Run}(e,p,t,b)$ (2.15)
to a single cubic instance representing the unbounded existential closure
$\exists t\;\mathsf{Run}(e,p,t,b).$ (2.16)
This is the point at which any putative uniform solver for $\mathcal{D}_{3}$ must supply a uniform witness-selection/packaging principle.
Definition 2.12 (Existential Closure of Bounded Runs).
For fixed $(e,p,b)$ define the $\Sigma_{1}$ predicate
$\mathsf{HaltsOut}(e,p,b)\;:\Longleftrightarrow\;\exists t\;\mathsf{Run}(e,p,t,b).$ (2.17)
Classically, $\mathsf{Solve}_{3}(e)$ asserts that for each input $p$ there exists a bit $b$ such that $\mathsf{HaltsOut}(e,p,b)$ holds, and that $b$ agrees with $\mathsf{Sol}_{3}(p)$.
Remark.
The bounded predicate $\mathsf{Run}(e,p,t,b)$ is decidable for fixed $t$. The unbounded closure $\exists t\,\mathsf{Run}(e,p,t,b)$ is exactly where non-uniformity enters: it asserts existence of a terminating trace without supplying a bound on where termination occurs. In realizability terms, it demands a witness-producing procedure, not merely a classical truth value.
Definition 2.13 (Cubic Closure Operator).
A cubic closure operator is a computable function $\mathsf{Close}$ which, given $(e,p,b)$, outputs a code
$q=\mathsf{Close}(e,p,b)\in\mathrm{Code}(\mathcal{D}_{3})$ (2.18)
such that
$\mathsf{HaltsOut}(e,p,b)\quad\Longleftrightarrow\quad\mathsf{Sol}_{3}\bigl(\mathsf{Close}(e,p,b)\bigr).$ (2.19)
Exposition.
One may equivalently regard $\mathsf{Close}$ as a uniform method for packaging a bounded-verification family (whose instance size depends on the bound $t$) into a single cubic Diophantine input while preserving solvability over $\mathbb{N}$. In this form, the closure problem is visibly orthogonal to the local “gadgetry” of Layer I: even perfect degree-$3$ bounded encodings do not by themselves provide a uniform elimination of the external bound.
Proposition 2.14 (Semantic Consequence).
Assume there exists $e$ such that $\mathsf{Solve}_{3}(e)$ holds. Then there exists a cubic closure operator $\mathsf{Close}$ satisfying (2.19).
Consider the predicate
$\mathsf{HaltsOut}(e,p,b)\equiv\exists t\,\mathsf{Run}(e,p,t,b)$ (2.20)
By Theorem 2.7, for each fixed $t$ there is a cubic system $\Phi_{e,p,t,b}=0$ whose solvability is equivalent to $\mathsf{Run}(e,p,t,b)$. The global solver hypothesis supplies a uniform terminating computation witnessing the correct yes/no decision for every cubic input. In particular, it supplies (uniformly in $(e,p,b)$) the computational content needed to compile the existential closure over $t$ into a single cubic solvability query: a code $q$ such that $q$ is solvable iff some $t$ yields a solvable bounded-run instance. Taking $\mathsf{Close}(e,p,b)\equiv q$ yields (2.19). ∎
Remark.
Proposition 2.14 is not an algebraic lemma about merging polynomials. It is a uniformity extraction: it isolates the implicit closure principle any total correct solver must realize in order to be globally correct on a class whose defining predicate already contains an unbounded existential quantifier. The remainder of the argument rejects the existence of such a closure operator at degree $3$ on realizability/non-uniformity grounds, and then applies a classical reflection step to conclude $\neg\exists e\,\mathsf{Solve}_{3}(e)$.
2.6 Constructive Reading
Exposition.
The closure principle isolated in Section 2.5 is not merely a classical equivalence between predicates.